OpenStudy (anonymous):

please help :) Use the arc length formula and the given information to find r. s = 16 cm, θ = 48°; r = ? sixty divided by pi cm thirty divided by pi cm one third cm one hundred twenty divided by pi cm

OpenStudy (anonymous):

\(s^2=\int_0^{48/360*2\pi}1+f'(x)^2dx\) putting in the values: \(16^2=\int_0^{48/360*2\pi}1dx+\int_0^{48/360*2\pi}f`(x)^2dx\) continue

OpenStudy (anonymous):

ok so that's the formula how to I cantinur to solve from there?

OpenStudy (anonymous):

continue

OpenStudy (debbieg):

Is this a trig class? are you using integrals, or the non-calculus arc length formula \[\Large s=\alpha r\] (with \(\alpha\) in radians)??

OpenStudy (anonymous):

its my pre calc class

OpenStudy (anonymous):

just havinga a total brain fart on this one

OpenStudy (anonymous):

these are my choices sixty divided by pi cm thirty divided by pi cm one third cm one hundred twenty divided by pi cm

OpenStudy (debbieg):

Well, if it's pre-calc, then I doubt you're doing integrals yet. The formula \(\Large s=\alpha r\) should look familiar though. But you have to convert to radians. \[\Large \alpha \cdot\frac{ \pi }{ 180 }\] will convert from degrees to radians. Then set up the equation with the formula, and solve it for r.

OpenStudy (anonymous):

convert 48 degrees to radians then?

OpenStudy (debbieg):

yes

OpenStudy (debbieg):

Do it EXACTLY, don't reach for your calculator. ;)

OpenStudy (anonymous):

if so that's \[\frac{ 4pie }{ 15 }\]

OpenStudy (debbieg):

Yes, that's it.

OpenStudy (debbieg):

\pi will give you \(\pi\) in the equation editor, btw. :)

OpenStudy (anonymous):

so\[16=\frac{ 4pie }{ 15 } ?\]

OpenStudy (debbieg):

don't forget r! that's what you're solving for!

OpenStudy (anonymous):

so where does r go in that eqAUTION?

OpenStudy (anonymous):

oops caps

OpenStudy (debbieg):

\(\Large 16=\dfrac{ 4\pi }{ 15 }\cdot r\)

OpenStudy (anonymous):

ok then uhmmm times r on both sides orrrr?

OpenStudy (debbieg):

this is the general equation for arc length: \(\Large s=\alpha r\) you're just plugging in the parts, and now solve for r.

OpenStudy (debbieg):

If you multiply by r on both sides, you'll have \(r^2\) on the right, and and r on the left. That is not useful, you want to ISOLATE r, so you need to deal with the \(\Large \dfrac{ 4\pi }{ 15 }\) coefficient. NOT the r.

OpenStudy (anonymous):

mmmmmmmm soooooo do I multiply reciperic or multiply bothtop and bottom by 15? sorry im trying I promise

OpenStudy (anonymous):

so its \[\frac{ 60 }{ \pi }\] correct?

OpenStudy (debbieg):

yes, the most direct way is just to multiply both sides by the reciprocal (since you have a rational coefficient and need to get it to the other side).

OpenStudy (debbieg):

multiplying "top and bottom" by 15 would not be useful. :) But, you could also do it in "2 steps", multiplying both SIDES by 15, and then dividing both sides by 4... that would work as well.

OpenStudy (anonymous):

so the answer isn't 60 over pie?

OpenStudy (debbieg):

lol.. yes, it is, that was my "yes" above. That's the correct answer. :) But do YOU understand WHY?

OpenStudy (anonymous):

yes I do now thank you so much !

OpenStudy (debbieg):

You're welcome! :) happy to help.